OpenStudy (anonymous):

How do you simplify the following, can you please show me a step by step process? Thank you.

OpenStudy (anonymous):

Simplify what?

OpenStudy (ragingsquirrel):

Be patient.

OpenStudy (anonymous):

\[(\sqrt{2}/2)\times (\sqrt{3}/2)+(-\sqrt{2}/2)\times(1/2)\]

OpenStudy (anonymous):